OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −1

OpenStudy (anonymous):

These are the choices: −4 4 −3 3

OpenStudy (ivettef365):

multiply the first equation by 3 multiply the second equation by -2

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

5x=20 x=4

OpenStudy (anonymous):

alright thx

OpenStudy (anonymous):

wait where did you get 5 from

OpenStudy (anonymous):

9x+6y=18 -4x-6y=-1

OpenStudy (anonymous):

-4x-6x=2 not -1 since -1 is mult by -2